A classical problem in additive number theory is to find the minimum size condition on a subset A of a finite field such that the nonzero elements of the finite field can be written as a sum of two products of elements of A.

In this talk we shall see that a higher dimensional perspective along with some elementary properties of character sums allow us to get better results than the ”one-dimensional” perspective previously allowed.